OpenStudy (anonymous):

the concentration C of a certain chemical in the bloodstream t hours after injection into muscle tissue is given by C=3t/27+t^3, t>=0. a) when is the concentration greatest? b)what is the greatest concentration? c)what is lowest concentration and when does the lowest concentration occur?

OpenStudy (anonymous):

i am confused, is it really \[C=\frac{3t}{27}+t^3\]?

OpenStudy (anonymous):

because this function grows without bound, so the bigger \(t\) is the bigger \(C\) is

OpenStudy (anonymous):

are you sure this is the right formula?

OpenStudy (anonymous):

C= (3t)/(27+t^3) when t>=0

OpenStudy (anonymous):

\[C= \frac{ 3t }{ 27+t^3 }\]

OpenStudy (anonymous):

To find the min/max of a function, find when the derivative is 0

OpenStudy (anonymous):

first derivative test?

OpenStudy (anonymous):

what min and max? u mean local min and local max? or absolute min and absolute max?

OpenStudy (anonymous):

Absolute. You need to find the absolute extrema.

OpenStudy (anonymous):

after i found the absolute extrema, would that be the greatest concentration?

OpenStudy (anonymous):

There's gotta be an absolute max and absolute min according to the question.

OpenStudy (anonymous):

how can i find the time of the greatest concentration?

OpenStudy (anonymous):

Find the first and second derivative. Then find the stationary points. \[C=\frac{ 3t}{ 27+t^3 }\] \[C'=\frac{3(27+t^3)-3t^2(3t)}{(27+t^3)^2}\] \[=\frac{81-6t^3}{(27+t^3)^2}\] \[C"=\frac{-18t^2(27+t^3)^2-6t^2(27+t^3)(81-6t^3)}{(27+t^3)^4}\] \[=\frac{-486t^2-18t^5-486t^2+36t^5}{(27+t^3)^3}\] \[=\frac{-972t^2+18t^5}{(27+t^3)^3}\] At S.P's(Stationary Points) C'=0 \[81=6t^3\] \[t^3=\frac{27}{2}\] \[t=\frac{3}{\sqrt[3]{2}}\] \[\huge C=\frac{ \frac { 9 }{ \sqrt[3]{2} } }{ 27+\frac{ 27 }{ 2 } } \] \[C=\frac{2}{9\sqrt[3]{2}}\]

OpenStudy (anonymous):

\[C''<0\] Therefore maximum.

OpenStudy (anonymous):

So I did a) and b) for you already.